Circular No. 6273

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1993J IN NGC 3031 M. J. Montes, N. E. Kassim, and K. W. Weiler, Naval Research Lab-

oratory; R. A. Sramek, National Radio Astronomy Observatory; and S. D. Van Dyk, University of California at Berkeley, report that new observations of SN 1993J made with the Very Large Array (VLA) in “B” configuration show that, through application of wide-field mapping algorithms, radio emission from the supernova is now detectable at wavelength 90 cm (0.330 GHz). On Oct. 6, the source had reached a level of 18.6 ± 2.9 mJy. SN 1993J remains strong and optically thin at higher frequencies, with flux densities on that date decreasing from 99.9 mJy at 1.4 GHz to 10.5 mJy at 22.5 GHz (spectral index of −0.83). Results at earlier times are given by Van Dyk et al. (1995, Ap.J. 432, L115). Monitoring with the VLA in all bands continues.

AX J2315–592 P. McGee, Department of Physics and Mathematical Physics, Uni-

versity of Adelaide, writes: “Differential CCD photometry of the optical counterpart of AX J2315–592 was undertaken at Woomera on Nov. 20–22 with a clear-filtered CCD. Exposures were 75 s long at 120-s intervals. Light curves covering somewhat over four cycles show a clear period as per IAUC 6260 and 6261 and exhibit different peak-trough amplitudes, as follows: Nov. 20.5035 UT (start of 1.5-hr run), amplitude 0.9 mag; 21.5378 (start of 2-hr run), 1.1 mag; 22.5041 (start of 3.7-hr run), 0.7 mag. Standard deviations for the comparison/check-star data were in the range 0.037–0.058 mag. It is expected that similar observations will be made over the period Dec. 12–22.”

COMET C/1995 O1 (HALE-BOPP) Corrigendum. On IAUC 6229, line 4, for K, 5.6 × 10^{−15} read K, 5.6

× 10^{−16}

Total magnitude estimates (cf. IAUC 6252): Oct. 28.11 UT, 10.0 (C. S. Morris, southern CA, 0.26-m reflector); Nov. 11.06, 10.1 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 14.06, 10.0 (Hale); 17.08, 9.6 (Morris); 19.08, 9.7 (Morris); 21.41, 10.0 (P. Camilleri, Cobram, Victoria, Australia, 0.20-m reflector); 24.07, 9.8: (Hale, 0.2-m reflector).

1995 December 12 Daniel W. E. Green